# Linear Bandits with Partially Observable Features

### v1

@@ 4 Robust Estimation for Partially Observable Features
We propose our estimation method to obtain a sublinear regret bound for linear bandits with latent features. Section 4.1 introduces the feature vector augmentation to handle the misspecification error and Section 4.2 presents the doubly robust estimation to further improve the regret bound.
@@ 4.1 Feature Vector Augmentation with Orthogonal Projection
In order to minimize regret, it is sufficient to estimate the $K$ expected rewards $\{\mathbf{z}_{a}^{\top}\boldsymbol{\theta}_{\star}:a\in[K]\}$ rather than all components of $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$. A straightforward approach to this problem, which achieves a regret bound of $\widetilde{O}(\sqrt{KT})$, is to disregard the observed features and apply MAB algorithms like UCB1 (Auer et al., 2002). However, these algorithms tend to incur higher regret than those leveraging features, particularly when the number of arms is significantly larger than the dimension of the feature vectors, i.e., $K\gg d$.
@@
We propose a unified approach to handle all cases of partially observable features and efficiently estimate all $K$ expected rewards. Let $\mathbf{X}:=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{d\times K}$ represent a matrix that concatenates the observed part of the true features, and $\mathbf{U}:=(\mathbf{u}_{1}^{(u)},\ldots,\mathbf{u}_{K}^{(u)})\in\mathbb{R}^{d_{u}\times K}$ represent the matrix that concatenates the latent complements of the true features for each arm. We assume a set of $K$ vectors $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ spans $\mathbb{R}^{d}$, without loss of generality. [^3] When $d>K$, we can apply singular value decomposition on $\mathbf{X}$ to reduce the feature dimension to $\bar{d}\leq K$ with $\mathrm{R}(\mathbf{X})=\bar{d}$. We define $\mathbf{P}_{\mathbf{X}}:=\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}$ as the projection matrix onto the row space of $\mathbf{X}$, denoted $\mathrm{R}(\mathbf{X})$. Then the vector of rewards for all arms, $\mathbf{Y}_{t}=(y_{1,t},\ldots,y_{K,t})$, is now decomposed as:
@@
$\mathbf{Y}_{t}$ $=(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+\epsilon_{t}\boldsymbol{1}_{K}$
$=\mathbf{P}_{\mathbf{X}}(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+\epsilon_{t}\boldsymbol{1}_{K}$
$=\mathbf{X}^{\top}(\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}+\epsilon_{t}\boldsymbol{1}_{K},$ (4)
@@
where the first and the second term are the projected rewards onto $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{X})^{\perp}$, the subspace of $\mathbb{R}^{K}$ perpendicular to $\mathrm{R}(\mathbf{X})$. We write the projected parameter as ${\boldsymbol{\mu}_{\star}^{(o)}}:=\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}$.
@@
To handle the second term in Eq. 4, we consider a set of basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}\in\mathrm{R}(\mathbf{X})^{\perp}$. Given the set, there exist coefficients ${\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}\in\mathbb{R}$ that express the reward projection as:
@@
$d_{h}(\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}):=|\{i\in[K-d]:\mu_{\star,i}^{(u)}\neq 0\}|.$ (6)
@@
Note that $d_{h}=0$ for any basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}$ when the latent feature space is included in the observed feature space, i.e., $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$, since it implies $(\mathbf{I}_{K}-\mathbf{P}_{X})\mathbf{U}^{\top}=\mathbf{0}_{K\times d_{u}}$. If $\mathrm{R}(\mathbf{U})\supseteq\mathrm{R}(\mathbf{X})$, then $d_{h}=K-d$ for any basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}$. In other cases, the quantity $d_{h}$ depends on the choice of the basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}$. Intuitively, $d_{h}$ tends to be small when $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$ have a large intersection, while some choice of the basis may lead to a higher $d_{h}$ and consequently to a higher regret bound. Nevertheless, for any choice of the basis, our algorithm achieves $\widetilde{O}(\sqrt{(d+d_{h})T})$ regret without prior knowledge of $d_{h}$, which is smaller than $\widetilde{O}(\sqrt{KT})$ regret bound achieved by MAB algorithms ignoring features.
@@
If we define $\boldsymbol{\mu}_{\star}$ as $[({\boldsymbol{\mu}_{\star}^{(o)}})^{\top},({\boldsymbol{\mu}_{\star}^{(u)}})^{\top}]^{\top}\in\mathbb{R}^{K}$, where $\boldsymbol{\mu}_{\star}^{(u)}=[{\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}]^{\top}$, then Eq. 4 becomes $\mathbf{Y}_{t}=[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\boldsymbol{1}_{K}$, implying that the reward for each $a\in[K]$ is:
$\begin{split}y_{a,t}&=\mathbf{e}_{a}^{\top}\mathbf{Y}\\ &=\mathbf{e}_{a}^{\top}[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\\ &=[{\mathbf{x}_{a}}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t},\end{split}$ (7)
@@
where $\mathbf{e}_{a}\in\mathbb{R}^{K}$ is a standard basis, with elements all zero except for $1$ in the $a$-th coordinate. With this modification, the rewards are now represented as a linear function of the augmented feature vectors, $\widetilde{\mathbf{x}}_{a}:=[\mathbf{x}_{a}^{\top}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]^{\top}\in\mathbb{R}^{K}$, without any misspecification error. A toy example illustrating our strategy is shown in Figure 1.
@@
Figure 1: Illustration comparing conventional linear bandit algorithms (left) and our approach (right) in estimating rewards of $K=3$ arms. Conventional algorithms use only observed features and find estimates within $\mathrm{R}(\mathbf{X})$ thus accumulating errors from unobserved features. However, our strategy projects the latent part of the reward onto the orthogonal complement of $\mathrm{R}(\mathbf{X})$, denoted by $\mathbf{b}_{1}^{\top}\widehat{\boldsymbol{\theta}}_{t}^{(u)}$, which enables reward estimation in $\mathbb{R}^{K}$. Note that $\hat{\boldsymbol{\theta}}_{t}$ is the estimator of the parameter for observed features.
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathrm{R}(\mathbf{X})$$\mathbf{q}_{2}$$\mathbf{X}^{\top}\hat{\boldsymbol{\theta}}_{t}$
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathbf{q}_{2}$($\mathrm{R}(\mathbf{X})$$\mathbf{b}_{1}^{\top}\widehat{\boldsymbol{\theta}}_{t}^{(u)}$$\mathbf{X}^{\top}\hat{\boldsymbol{\theta}}_{t}$
@@
The dimension of the augmented feature vectors $\{\widetilde{\mathbf{x}}_{a}:a\in[K]\}$ is $K\geq d$. While applying SupLinUCB (Chu et al., 2011) yields $\widetilde{O}(\sqrt{KT})$ regret bound, it lacks practicality as it computes $\log T$ distinct batches and estimators requiring the knowledge of $T$ and, more critically, discards a significant portion of samples in each parameter update. We present an efficient algorithm that employs the doubly robust ridge estimator and achieves $\widetilde{O}(\sqrt{KT})$ regret bound (see Appendix A). However, when $K>d$ and $d_{u}=0$, the regret is high compared to the linear bandits with conventional features. Therefore, we propose a novel estimation strategy to avoid dependency on $K$ in the following section.
@@ 4.2 Doubly Robust Lasso Estimator
In Eq. 7, the parameter $\boldsymbol{\mu}_{\star}$ is sparse depending on the dimension of the latent features and the augmented basis vectors. Recall that $\boldsymbol{\mu}_{\star}^{(u)}$ are the coefficients to express the projection of the reward as represented in Eq. 5 and only $d_{h}$ basis vectors are required to express the projection of the reward; there are at most $d_{h}$ nonzero entries in $\boldsymbol{\mu}_{\star}^{(u)}$.
@@
Let $\check{\boldsymbol{\mu}}^{L}_{t}$ denote the Lasso estimator for $\boldsymbol{\mu}_{\star}$ using augmented feature vectors:
@@
$\begin{split}\check{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}(y_{a_{\tau}}-\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}\boldsymbol{\mu})^{2}+2\sigma\sqrt{\frac{2t}{p}\log\frac{2Kt^{2}}{\delta}}\Big{\|}\big{(}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}\big{)}^{\frac{1}{2}}\boldsymbol{\mu}\Big{\|}_{1}.\end{split}$ (8)
@@
For the estimator in Eq. 8 to correctly identify the zero entries in $\boldsymbol{\mu}_{\star}$, the compatibility condition must hold (van de Geer and Bühlmann, 2009), which requires the minimum eigenvalue of the Gram matrix, $\lambda_{\min}\left(t^{-1}\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top}\right)$, to be positive. However, ensuring a sufficiently large minimum eigenvalue typically demands collecting a large number of exploration samples, thus increasing regret. Achieving this with fewer exploration samples remains a key challenge in bandit literature, as the minimum eigenvalue influences convergence rate of the estimator and, consequently, the regret bound (Kim et al., 2021; Soare et al., 2014).
@@
We introduce a doubly robust (DR) estimator that employs the full feature Gram matrix $\sum_{s=1}^{t}\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$ instead of $\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top}$. The DR estimation originates from the statistical literature on missing data, where “doubly robust” means that the estimator is robust against errors in the estimation of both the observation probability and the response model. In bandits, at each decision round $t\in[T]$, only the reward of the selected arm is observed, while the $K-1$ unselected rewards are missing. Thus DR estimation is applied to impute these $K-1$ missing rewards and include corresponding $K-1$ feature vectors in the estimation. Since the observation probability is given by the policy (which is known to the learner), the DR estimator is robust against errors in the estimated rewards. While Kim and Paik (2019) proposed a DR Lasso estimator on IID features satisfying the compatibility condition, we propose another DR Lasso estimator that does not require the assumptions on features.
@@
We improve the DR estimation by incorporating resampling and coupling methods. For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote an exploration phase such that for $\tau\in\mathcal{E}_{t}$ the action $a_{\tau}$ is sampled uniformly over $[K]$. The $\mathcal{E}_{t}$ is constructed as $\mathcal{E}_{0}=\emptyset$ and $\mathcal{E}_{t}=\mathcal{E}_{t-1}\cup{t}$ if $|\mathcal{E}_{t}|\leq C_{e}\log(2Kt^{2}/\delta)$, otherwise $\mathcal{E}_{t}=\mathcal{E}_{t-1}$, where $C_{e}:=8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$. In round $t$, when $t\notin\mathcal{E}_{t}$, the algorithm selects an action $a_{t}$ according to an $\epsilon_{t}$-greedy policy. Then, we generate a pseudo-action $\widetilde{a}_{t}$ from a multinomial distribution:
$\begin{split}&phi_{a_{t},t}:=\mathbb{P}(\widetilde{a}_{t}=a_{t}|a_{t})=p\quad\text{and}\quad\phi_{k,t}:=\mathbb{P}(\widetilde{a}_{t}=k|a_{t})=\frac{1-p}{K-1}\quad\forall k\in[K]\setminus\{a_{t}\},\end{split}$ (9)

### v2

@@ 4.1 Feature Vector Augmentation with Orthogonal Projection
In the linear bandit framework, accurate estimation of the unknown reward parameter $\boldsymbol{\theta}_{\star}$ contributes to optimal decision making. However, in our problem setting, the learner lacks access to the latent reward component—namely, the last term in Eq. 3. Consequently, without any specific treatment to compensate for the absence of $\boldsymbol{\theta}_{\star}^{(u)}$, the agent fails to accurately estimate $\boldsymbol{\theta}_{\star}$, resulting in ineffective learning, as demonstrated in in Theorem 1.
@@
That said, minimizing regret—that is, selecting the optimal arm—does not require recovering the true reward parameter $\boldsymbol{\theta}_{\star}$ entirely; rather, it suffices to estimate the $K$ expected rewards $\{\mathbf{z}_{a}^{\top}\boldsymbol{\theta}_{\star}:a\in[K]\}$. A straightforward approach is to ignore the features altogether and apply MAB algorithms such as UCB1 (Auer et al., 2002), which are known to achieve a regret bound of $\widetilde{O}(\sqrt{KT})$. However, these algorithms tend to incur higher regret than those leveraging features, particularly when the number of arms is significantly larger than the dimension of the feature vectors, i.e., $K\gg d_{z}$.
@@
To tackle this dilemma, we propose a unified approach to handle all cases of partially observable features and efficiently estimate the expected rewards. Let $\mathbf{X}:=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{d\times K}$ represent a matrix that concatenates the observed part of the true features, and $\mathbf{U}:=(\mathbf{u}_{1}^{(u)},\ldots,\mathbf{u}_{K}^{(u)})\in\mathbb{R}^{d_{u}\times K}$ represent the matrix that concatenates the latent complements of the true features for each arm. We assume a set of $K$ vectors $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ spans $\mathbb{R}^{d}$, without loss of generality. [^3] When $d>K$, we can apply singular value decomposition on $\mathbf{X}$ to reduce the feature dimension to $\bar{d}\leq K$ with $\mathrm{R}(\mathbf{X})=\bar{d}$. We define $\mathbf{P}_{\mathbf{X}}:=\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}$ as the projection matrix onto the row space of $\mathbf{X}$, denoted by $\mathrm{R}(\mathbf{X})$. Then the vector of rewards for all arms, $\mathbf{Y}_{t}=(y_{1,t},\ldots,y_{K,t})^{\top}$, is now decomposed as:
$\mathbf{Y}_{t}$ $=\left(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\right)+\epsilon_{t}\boldsymbol{1}_{K}$
$=\mathbf{P}_{\mathbf{X}}\left(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\right)$
@@
$\quad+(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})\left(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\right)+\epsilon_{t}\boldsymbol{1}_{K}$
$=\mathbf{X}^{\top}\left(\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\right)$
$\quad+(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}+\epsilon_{t}\boldsymbol{1}_{K},$ (4)
@@
where the first term corresponds to the reward projected onto $\mathrm{R}(\mathbf{X})$; while the second term is the projected rewards onto $\mathrm{R}(\mathbf{X})^{\perp}$, the subspace of $\mathbb{R}^{K}$ perpendicular to $\mathrm{R}(\mathbf{X})$. We write the projected parameter as ${\boldsymbol{\mu}_{\star}^{(o)}}:=\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}$.
@@
To handle the second term in Eq. 4, we consider a set of row vector basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}\in\mathrm{R}(\mathbf{X})^{\perp}$, where $\mathbf{b}_{i}\in\mathbb{R}^{K}$ for $i\in[K-d]$. Given the set, there exist coefficients ${\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}\in\mathbb{R}$ to express the reward projection as:
@@
$d_{h}(\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}):=|\{i\in[K-d]:\mu_{\star,i}^{(u)}\neq 0\}|.$ (6)
Note that $d_{h}=0$ for any basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$ when the latent feature space is completely included in the observed feature space, i.e., $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$. In this case, $(\mathbf{I}_{K}-\mathbf{P}_{X})\mathbf{U}^{\top}=\mathbf{0}_{K\times d_{u}}$, which means that the projected rewards onto $\mathrm{R}(\mathbf{X})^{\perp}$ can be linearly expressed by the observed features. If $\mathrm{R}(\mathbf{U})\supseteq\mathrm{R}(\mathbf{X})$, on the other hand, then $d_{h}=K-d$ for any basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$.
@@
In the other cases, the quantity $d_{h}$ depends on the choice of the basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$, and tends to be smaller when a larger portion of $\mathrm{R}(\mathbf{U})$ can be expressed within $\mathrm{R}(\mathbf{X})$. For any choice of the basis, our algorithm achieves $\widetilde{O}(\sqrt{(d+d_{h})T})$ regret without prior knowledge of $d_{h}$, which does not exceed $\widetilde{O}(\sqrt{KT})$ regret bound achieved by MAB algorithms ignoring features. Further details are provided in Section 5.2.
@@
Let us define $\boldsymbol{\mu}_{\star}$ as $[({\boldsymbol{\mu}_{\star}^{(o)}})^{\top},({\boldsymbol{\mu}_{\star}^{(u)}})^{\top}]^{\top}\in\mathbb{R}^{K}$, where $\boldsymbol{\mu}_{\star}^{(u)}=[{\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}]^{\top}$, then the reward for each $a\in[K]$ can be modified as follows:
$\begin{split}y_{a,t}&=\mathbf{e}_{a}^{\top}\mathbf{Y}\\ &=\mathbf{e}_{a}^{\top}[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\\ &=[{\mathbf{x}_{a}}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}.\end{split}$ (7)
@@
The decomposition in Eq. 7 implies that Eq. 4 takes the form $\mathbf{Y}_{t}=[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\boldsymbol{1}_{K}$. Note that $\mathbf{e}_{a}\in\mathbb{R}^{K}$ is a Euclidean basis, with elements all zero except for $1$ in the $a$-th coordinate. With this modification, the rewards are now represented as a linear function of the augmented feature vectors: $\widetilde{\mathbf{x}}_{a}:=[\mathbf{x}_{a}^{\top}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]^{\top}\in\mathbb{R}^{K}$, without any misspecification error. A toy example illustrating our strategy is shown in Figure 1.
@@
Figure 1: Illustration comparing OLB algorithms and our approach in estimating rewards of $K=3$ arms. OLB algorithms find estimates within $\mathrm{R}(\mathbf{X})$ thus accumulating errors from unobserved features. However, our approach utilizes projection of the latent reward, $\mathbf{b}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(u)}$, onto the orthogonal complement of $\mathrm{R}(\mathbf{X})$, enabling reward estimation in $\mathbb{R}^{K}$. Note that $\widehat{\boldsymbol{\mu}}_{t}$ is the estimator of the parameter for observed features.
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathrm{R}(\mathbf{X})$$\mathbf{q}_{2}$$\mathbf{X}^{\top}\widehat{\boldsymbol{\mu}}_{t}$
(a) OLB algorithms: Estimation confined to $\mathrm{R}(\mathbf{X})$
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathbf{q}_{2}$($\mathrm{R}(\mathbf{X})$$\mathbf{b}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(u)}$$\mathbf{X}^{\top}\widehat{\boldsymbol{\mu}}_{t}$
(b) Ours: Projection onto orthogonal complement
@@
The dimension of the augmented feature vectors $\{\widetilde{\mathbf{x}}_{a}:a\in[K]\}$ is $K\geq d$. While applying SupLinUCB (Chu et al., 2011) yields $\widetilde{O}(\sqrt{dT})$ regret bound, it is known to lack practicality as it computes $\log T$ distinct batches and estimators requiring the knowledge of $T$ and, more critically, discards a significant portion of samples in each parameter update. We present an efficient algorithm that employs the doubly robust ridge estimator and achieves $\widetilde{O}(\sqrt{KT})$ regret bound (see Appendix C). However, when $K>d$ and $d_{u}=0$, the regret is high compared to the linear bandits with conventional features. To address this challenge, we propose a novel estimation strategy in the following section that eliminates the dependence on $K$ in the regret bound.
@@ 4.2 Doubly Robust Lasso Estimator
In Eq. 7, the parameter $\boldsymbol{\mu}_{\star}$ is sparse, with its sparsity depending on the dimension of the subspace that captures the inaccessible portion of the reward—$d_{h}$ defined in Eq. 6. Recall that $\boldsymbol{\mu}_{\star}^{(u)}$ contains the coefficients used to express the projection of the reward, as represented in Eq. 5, and that only $d_{h}$ basis vectors are required to express this projection. Therefore, $\boldsymbol{\mu}_{\star}^{(u)}$ contains at most $d_{h}$ nonzero entries.
@@
Let $\check{\boldsymbol{\mu}}^{L}_{t}$ denote the Lasso estimator of $\boldsymbol{\mu}_{\star}$ using the augmented feature vectors, defined as follows:
@@
$\begin{split}\check{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}(y_{a_{\tau},\tau}-\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}\boldsymbol{\mu})^{2}\\ &\qquad+2\widetilde{\sigma}_{\max}\sigma\sqrt{2pt\log\frac{2Kt^{2}}{\delta}}\|\boldsymbol{\mu}\|_{1},\end{split}$ (8)
where $p$ is the coupling probability used to define the multinomial distribution for pseudo-action sampling (as defined in Eq. 9), and $\delta$ is the confidence parameter of the algorithm. Note that $\widetilde{\sigma}^{2}_{\max}:=\max_{a\in[K]}\mathbf{e}_{a}^{\top}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\mathbf{e}_{a}$, i.e., the largest diagonal entry of the Gram matrix constructed from the augmented feature vectors $\widetilde{\mathbf{x}}_{a}$ over $\mathcal{A}$.
@@
For the estimator in Eq. 8 to correctly identify the zero entries in $\boldsymbol{\mu}_{\star}$, the compatibility condition is required to be satisfied (van de Geer & Bühlmann, 2009). Although the compatibility condition does not, in general, require a positive minimum eigenvalue, in our setting the Gram matrix $\lambda_{\min}(t^{-1}\sum_{s=1}^{t}\widetilde{x}_{a_{s}}\widetilde{x}_{a_{s}}^{\top})$ has a strictly positive minimum eigenvalue. Therefore, the compatibility condition is implied without any additional assumption. However, ensuring a sufficiently large minimum eigenvalue typically requires collecting a large number of exploration samples, which in turn increases regret. Achieving this with fewer exploration samples remains a key challenge in bandit literature, as the minimum eigenvalue affects the convergence rate of the estimator and, consequently, the regret bound (Soare et al., 2014; Kim et al., 2021).
@@
We introduce a doubly robust (DR) estimator that employs the Gram matrix constructed from the augmented feature vectors over the entire action space, $\sum_{s=1}^{t}\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$, instead of the Gram matrix based only on the vectors chosen at each round, i.e., $\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top}$. The DR estimation originates from the statistical literature on missing data (Bang & Robins, 2005), where “doubly robust” refers to estimators that remain consistent if either the imputation model or the observation probability model is correctly specified (Kim et al., 2021). In the bandit setting, at each decision round $t\in[T]$, only the reward corresponding to the selected arm is observed, while the $K-1$ unselected rewards are missing. Hence, DR estimation imputes these $K-1$ missing rewards and incorporates the associated feature vectors into the estimation process. Since the observation probabilities are determined by the policy—which is known to the learner—the DR estimator remains robust to errors in the reward estimation. While Kim & Paik (2019) proposed a DR Lasso estimator under IID features satisfying the compatibility condition, we propose a novel DR Lasso estimator that does not rely on such assumptions.
@@
We improve the DR estimation by incorporating resampling and coupling methods. For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote the set of exploration rounds such that for any $\tau\in\mathcal{E}_{t}$, the action $a_{\tau}$ is sampled uniformly over $[K]$. The set $\mathcal{E}_{t}$ is constructed recursively. Starting with $\mathcal{E}_{0}=\emptyset$, we define
$\mathcal{E}_{t}=\begin{cases}\mathcal{E}_{t-1}\cup\{t\}&\mathrm{if}\quad|\mathcal{E}_{t}|\leq C_{e}\log\dfrac{2Kt^{2}}{\delta},\\ \mathcal{E}_{t-1}&\mathrm{otherwise}.\end{cases}\vspace{-0.2cm}$
Here, $C_{e}$ is defined as $(8K)^{3}\widetilde{\sigma}_{\min}^{-2}\widetilde{\sigma}_{\max}^{2}(1-p)^{-2}$, where $\widetilde{\sigma}_{\min}^{2}:=\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})$ denotes the minimum eigenvalue of the augmented Gram matrix over $\mathcal{A}$.
At each round $t$, when $t\notin\mathcal{E}_{t}$, the algorithm selects an action $a_{t}$ according to an $\epsilon_{t}$-greedy policy. Then, a pseudo-action $\widetilde{a}_{t}$ is generated from a multinomial distribution:
$\begin{split}&\phi_{a_{t},t}:=\mathbb{P}(\widetilde{a}_{t}=a_{t}\mid a_{t})=p,\\ &\phi_{k,t}:=\mathbb{P}(\widetilde{a}_{t}=k\mid a_{t})=\frac{1-p}{K-1},\end{split}$ (9)
for all $k\in[K]\setminus\{a_{t}\}$, where $p\in(1/2,1)$ is the coupling probability specified by the algorithm. To couple the policies for the actual action $a_{t}$ and the pseudo-action $\widetilde{a}_{t}$, we repeatedly resample both of them until they match.
When the DR estimation is applied under the $\epsilon_{t}$-greedy policy, the pseudo-reward (Eq. 10) involves inverse probability term $\epsilon_{t}^{-1}:=\sqrt{t}$, which can cause its variance to grow unbounded over time. To mitigate this issue, we propose a resampling and coupling strategy: by coupling the $\epsilon_{t}$-greedy policy with the multinomial distribution Eq. 9, we ensure that each inverse probability weight $\phi_{a,t}^{-1}$, for $a\in[K]$, remains bounded by $O(K)$. This coupling yields a lower bound on the observation probability, thereby reducing the variance of the DR pseudo-rewards (Eq. 10).